SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 20 August 2005

Status Report From: NASA HQ/SpaceRef/NASA Watch
Posted: Saturday, August 20, 2005

image SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  Saturday -- off-duty day for Sergei Krikalev and John Phillips, except for some housekeeping, cleanup and voluntary work. 

Crew sleep cycle is back on regular schedule, with wake-up at 2:00am EDT and sleep begin at 5:30pm.

As part of today's morning inspection after wakeup, CDR Krikalev did the periodic checkup behind panel 139 in the Service Module (SM) on a fluid connector of the urine collection system, checking for potential moisture.

Today's final post-EVA cleanup activities by Krikalev consisted of removing oxygen tanks (BK3), telemetry systems (BRTA) and 825-3M batteries from the Orlan-M suit backpacks, setting the first of the two 825-3Ms up for discharging, refilling the spacesuits' feedwater tanks with water, and arranging the Orlans to dry out during the day's course.

Later tonight, Sergei will stow the dried-out Orlan suits and their BSS interface control units. 

The CDR also completed closeout operations on the retrieved Japanese MPAC/SEED payload containers for secure stowage and subsequent return to Earth.

Shuttle
Processing Status
News
Daily Mission
Return to Flight
ISS
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

FE Phillips meanwhile stowed the tools used during the EVA.  Afterwards, John returned the U.S. SODF (Systems Operation Data File), Emergency books, CCPK (Crew Contamination Protection Kit) plus CD Library from their temporary location in the DC1 docking module to their regular places in the Lab.

Sergei completed the regular daily maintenance of the SM ECLSS (SOZh) systems plus ASU toilet facilities. 

Both crewmembers conducted their regular 2.5-hr. physical exercise program on the TVIS treadmill, CEVIS cycle ergometer, RED resistive machine and VELO bike with bungee cord load trainer.   [Sergei's daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO plus load trainer (today: Day 1 of a new set).]

At ~3:40am, Krikalev held a private conference with friends at TsUP/Moscow via S- & Ku-band.

At ~11:20am, the crew conducted their regular weekly planning conference (WPC) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners), via S-band/audio, reviewing upcoming activities and any concerns about future on-orbit events.

At ~11:40am, Phillips had his weekly private conference with his family (PFC), via S-band/audio and Ku-band/NetMeeting video (which employs an USB camera at the SSC laptop).

Station attitude is now XPOP (x-axis perpendicular to orbit plane), and the P6 solar array wings are set to dual-angle autotrack, with BGA (Beta gimbal assembly) 2B at -36 deg, 4B at +36 deg angle.

Vozdukh is still down, CDRA (CO2 removal assembly) is running in single-bed mode, and the TCCS (trace contaminants control subsystem) is off.

Correction to today's ISS On-Orbit Status:

Vozdukh was reactivated last night and continues to operate in manual mode 5.   CDRA continues to operate also, currently in dual-bed mode.   [Early this morning, three good pumpo-downs were completed in single-bed mode, indicating that the failed check valve has re-seated.  MCC-Houston and TsUP/Moscow are considering whether to deactivate CDRA later tonight following a minimum of one good cycle in dual-bed operations.  CO2 levels

Weekly Science Update (Expedition Eleven -- 17th): 

Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology (HRF GASMAP):  After the recent GASMAP "health" check, payload specialists expect that it was the last in HRF Rack 1, with the next one after GASMAP's move to the newly arrived HRF Rack 2. 

Human Research Facility/Workstation (HRF WS):  Continuing. 

Advanced Ultrasound (ADUM):  Continuing.

Renal Stone (RS):  In progress.

Foot/Ground Reaction Forces during Space Flight (FOOT):   Complete.

Space Acceleration Measurement System (SAMS):  SAMS is Off. 

Microgravity Acceleration Measurement System (MAMS):   MAMS remains in nominal operations. 

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): Complete and returned to the PI. 

Binary Colloidal Alloy Test-3 (BCAT-3):    BCAT-3 Slow Growth Sample Module will be left undisturbed in its current location by the E11 crew.  In order for the samples to potentially grow crystals that can be photographed during Increment 12 operations, the Sample Module must be left undisturbed. 

Materials ISS Experiment (MISSE):  In progress.  New MISSE-5 "suitcase" deployed and unfolded during EVA outside on the U.S. Airlock.  Reactivated after EVA-14.  Nominal and collecting data.  Old unit was returned on LF-1.

Dust and Aerosol Measurement Feasibility Test (DAFT):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):   Complete.

Serial Network Flow Monitor (SNFM):  Complete.

Fluid Merging Viscosity Measurement (FMVM):  Nothing new.

Space Experiment Module (SEM):  Nothing new.  Experimenters and kids are working to get the next two satchels on ULF1.1. 

Miscible Fluids in Microgravity (MFMG):   MFMG payload operations are finished. 

Educational Payload Operations (EPO):   Complete. 

Crew Earth Observations (CEO):   Through 8/14, a total of 4,984 CEO images were received on the ground.  The crew's continued practice with the long lens setting is reportedly paying off, and investigators are very pleased with the focus quality of many of the recent images.  However, the return to XPOP attitude and the Southern Hemisphere for daylight-awake periods will reduce CEO opportunities for more for awhile.  An extraordinary image of Mt. McKinley will be posted as Image of the Day on NASA/GSFC's Earth Observatory website next week.  Besides providing an unprecedented, detailed, high-oblique view of North America's highest peak from Space, this photo also documents part of an enormous forest fire smoke pall covering much of Alaska. 

Today's optional CEO photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), were Internal waves, Azores, Atlantic (on this fair weather pass the crew was to look well forward and left of track to find the Azores themselves in the area of sun glint and begin shooting.  Continuing for about another minute as the glint area moves southeastward), California Current Ecosystem (light was almost gone for the area, but the weather was fair.  Looking just left of track for southern California's Channel Islands and trying to keep the islands in the field of view and map the waters to the SW and S), and Internal waves, N Patagonian Shelf (the weather south of the Valdes Peninsula may just hold off for one more day.  After ISS crossed the coast, the crew was to begin looking for sea surface features in the glint near and southeast of the Valdes).

CEO photography can be viewed and studied at the websites:

See also the website "Space Station Challenge" at:

To view the latest photos taken by the expedition 11 crew visit:

Expedition 11 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS Tracker - More Links

ISS Orbit  (as of this morning, 11:34am EDT [= epoch]):

  • Mean altitude -- 353.6 km
  • Apogee height -- 354.4 km
  • Perigee height -- 352.8 km
  • Period -- 91.61 min.
  • Inclination (to Equator) -- 51.64 deg
  • Eccentricity -- 0.0001191
  • Solar Beta Angle -- 36.0 deg (magnitude increasing)
  • Orbits per 24-hr. day -- 15.72
  • Mean altitude loss in last 24 hours -- 110 m
  • Revolutions since FGB/Zarya launch (Nov. 98) -- 38582


ISS Altitude History

Apogee height -- Mean Altitude -- Perigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter